Phenol, C6H5OH, has a Ka of 1.3×10−10.

1.Write out the Ka reaction for phenol.

2. Calculate Kb for phenol’s conjugate base.

3.Is phenol a stronger or weaker acid than water? explain why.


1
Expert's answer
2021-02-28T06:15:24-0500

a. C6H5OH + H2O ---> C6H5O- + H3O+


b.Kb(C6H5O−)=Kw/Ka(C6H5OH)

Kb(C6H5O−)=Kw/Ka(C6H5OH)

Calculate the value of the base dissociation constant of phenoxide ion by substituting the given value of the acid dissociation constant of phenol and thatof ionic product of water as follows:

Kb(C6H5O−)=1.0×10-14/1.00×10-10

=1.00×10-4


c. phenol is more acidic than water too, because water has more polar OH-group (in H-OH) than in phenol, because, the alkyl group releases electrons and minimizes polarity of -OH group so, the water can have more stable hydroxide ion.